OpenStudy (anonymous):

Solve 4x2 = 20x. Answer x = 4, x = 5 x = 4, x = -5 x = 0, x = 5 x = 0, x = -5

OpenStudy (anonymous):

correct.. is... 4x^2 = 20x

OpenStudy (aravindg):

5

OpenStudy (aravindg):

cancel off x 4x=20 x=5

OpenStudy (aravindg):

also not expression holds for x=0 too

OpenStudy (aravindg):

so x=5,x=0

OpenStudy (anonymous):

x = 0, x = 5

jimthompson5910 (jim_thompson5910):

4x^2 = 20x 4x^2 - 20x = 0 4x(x-5) = 0 4x = 0 or x-5 = 0 x = 0 or x = 5

OpenStudy (aravindg):

note that @jim_thompson5910 used the systematic method for doing sch qns , what i did is also right bt before cancelling the x on both sides i looked at the condition when x=0

OpenStudy (anonymous):

ok thanks guys shows me two diffrent ways to do it..

OpenStudy (aravindg):

it is adviced u clearly understand our wrking